OpenStudy (anonymous):

does 1/2sin^2x= -1/4 cos(2x)

OpenStudy (anonymous):

Expand out both sides to see if that's true

OpenStudy (anonymous):

not sure if I know how to do that..

OpenStudy (anonymous):

cos(2x)=cos(x+x). now tell me how to expand the right hand side?

OpenStudy (anonymous):

2sin(x)^2=-cos(2x) cos(2X)=1-2sin^2x so.. 2sin(x)^2=2sin(x)^2-1 ?

OpenStudy (anonymous):

yeah, actually. you're done!

OpenStudy (anonymous):

so not equal... :(

OpenStudy (anonymous):

does

OpenStudy (anonymous):

yeah, you arrived at a contradiction and it can't be true.

OpenStudy (anonymous):

\[\int\limits_{?}^{?} cosxsinx dx\] = which one?

OpenStudy (anonymous):

integral of cosx * sinx dx I got the 1/2 sin^2(x) but my answer key I have the other one

OpenStudy (anonymous):



OpenStudy (anonymous):